Calculate the internal energy of one mole of nitrogen gas at pressure 50k pa occupying at V=5.10cm
1
Expert's answer
2021-10-24T18:30:02-0400

"i=5,"

"U=\frac i2 \nu RT=\frac i2pV,"

"U=\frac 52\cdot 5\cdot 10^4\cdot 5.1\cdot 10^{-6}=0.64~J."